How many grams of beryllium BeCl 2 are needed to make 125ml of 10 % solutions


1
Expert's answer
2021-02-05T04:04:40-0500

Solution:

Concentration in % (w/v) = 100 × (mass solute in g) / (volume solution in mL)

Hence,

10% m/v = 100 × (mass BeCl2 / 125 mL)

mass BeCl2 = 12.5 g 


Answer: 12.5 grams of BeCl2 are needed to make 125 mL of 10% solution.